Status Report

NASA Space Station On-Orbit Status 21 Feb 2004

By SpaceRef Editor
February 21, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 21 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.   As of this morning at ~6:39am EST, the ISS has circled Earth 30.000 times since FGB/Zarya launch.  Hurrah!

Flight Control to Crew: “We hope you are able to relax and enjoy your well-deserved weekend.  Happy Saturday!”

After an extended sleep-in (to 5:45am EST), Foale and Kaleri performed the regular 3-hr. Saturday task of uborka stantsii (station cleaning).

Mike tended the PromISS-3 experiment (Protein Crystal Growth Monitoring by Digital Holographic Microscope 3) in the smoothly running Microgravity Science Glovebox, today removing tape #10 and installing tape #11.

Alex Kaleri conducted the daily routine maintenance on the SOZh life support system, including the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO).  His maintenance/inspection tour also covered the weekly BRPK air/water condensate separator.

Sasha prepared the IMS (inventory management system) “delta” file required for updating the IMS databases.

The crew worked out according to their regular daily physical exercise program on TVIS treadmill and CEVIS cycle (aerobic), VELO ergometer with force loader and RED (anaerobic).   [According to Russian specialists from IBMP the VELO frame temperatures reported on 2/15 by the crew as high (70 degC est.), were expected.  There is no concern with the hardware, and no further investigation will occur.] 

At 9:10am, the crew conducted their weekly planning conference with the ground via S-band to discuss next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners and uplinked ahead of time).

At 1:30pm, they also had the monthly teleconference with ISS Program Management via S-band/audio.

Both crewmembers enjoyed their weekly private family conference (PFC), Mike via S- and Ku-band/video, Sasha via VHF/audio (Russian: UKV  ultra-short wave).

During yesterday’s instrument calibration, the Service Module gas analyzer (SM GA) and U.S. major constituents analyzer (MCA) were found to be within 1 mm/Hg (torr), an excellent agreement confirming their current accuracy.

The CDR was requested to take an additional ppCO2 reading in the SM with the CDM (carbon dioxide monitoring kit).   [These weekly readings in both the Lab and SM using the same CDMK unit will be conducted for the next few weeks to help the ground better assessif IMV (intermodular ventilation) flow between the U.S. and Russian segments has become degraded.]

At 1:40pm, ISS attitude control was handed over to the RS for Russian thrusters to change flight attitude from earth-oriented LVLH/YVV attitude to sun-pointing XPOP (x-axis perpendicular to orbit plane).  Return to LVLH is scheduled on 2/25, in time to support the EVA-9.   [The maneuver started at 2:11pm and lasted eight minutes.  Attitude control was returned to U.S. MCS (motion control system) and CMGs (control moment gyros) at 2:30pm for momentum management.]

A new pre-EVA activity added to Sasha’s Russian task list concerns requires him to clean his Nikon D1 digital camera’s highly sensitive CCD (charged-couple device) matrix of dust particles.   [Earlier in the program, when downlinked frames of the camera, which uses the AF-S Nikkor lens, had shown smudges looking like “soft focus spots”, Yuri Malenchenko had investigated possible causes and methods for cleaning the matrix. Kaleri is to determine if any contamination is present and, if so, whether he is able to remove it by softly blowing it off with air.]

After yesterday’s swap-out of Russian laptop 1 with the medical & payload laptop 3, TsUP/Moscow today (7:40am) used S-band to command SM Central Post computer #2 (KTsP2) off and switch operations back to KTsP1.   [Laptop 1, working with KTsP1, failed a week ago (2/12).  After operations were switched to the redundant (backup) laptop 2, with KTsP2, Kaleri replaced its HDD (hard disk drive) #6136 on 2/13 with a spare (#6059), but it again failed to reboot.  After yesterday’s laptop swap, Laptop 3’s previous medical and payload experiments can no longer be operated, until #1 has been repaired.  Laptop 3 is apparently is not required to support medical activities prior to or following the EVA-9.]

At dinnertime (7:15pm), as every day, the crew will support the Renal Stone prevention experiment by taking the test medication (either potassium citrate or placebo tablets) until the next sample collection phase in early April this year.

Tonight at ~8:00 pm, TsUP/Moscow plans to perform a calibration session on the critical BILU linear accelerometer unit in the Soyuz TMA-3/7S vehicle, using RGS (Russian ground stations) for command uplink and U.S. S-band for telemetry downlink (as for the preceding attitude change).

ECLSS specialists at MCC-Houston have prepared calculations for different repress options with Progress 13P air and/or oxygen before and after the EVA-9, depending on the operation of the Elektron O2 generator.   [A nonfunctioning Elektron will not put planned activities at risk.]

Yesterday’s condensate water offloading from the Lab collection tank by Foale transferred 57 lbs to a CWC (contingency water container).   [At the time of the offloading the tank was at 71 lbs.  The offload allows the tank to collect more condensate in the event that the Russian SKV-1 dehumidifier fails during the EVA-9 activities on 2/26.]

The crew’s sleep time begins tonight at 9:15pm EST.

Weekly Science Update (Expedition Eight — 16th):

GASMAP:   Next GASMAP operations will occur sometime after EVA-9.  The ground looks forward to working with the crew in early March

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound:   Planned.

Hand Posture Analyzer (HPA:  Looking forward to future operations.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Looking forward to future operations, after the PromISS operations.
Space Acceleration Measurement System (SAMS):   SAMS will capture the disturbances from the Orlan EVA on 2/26.
Microgravity Acceleration Measurement System (MAMS):   MAMS will captured the ISS maneuver to XPOP today and will also capture the maneuver back to LVLH on 2/25.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  The PROMISS 3 experiment has been running flawlessly for 19.5 days. The protein crystal growth environment has been very stable at 25degC +/- 0.3degC.  Experimental conditions are optimal and the optical microscope shows very stable fringe patterns.  On 2/16, due to the unexpected short length of tape #8, PROMISS 3 switched to WAIT TAPE mode earlier than foreseen; the situation reverted to a normal RUN mode after the tape #8 to tape #9 exchange.  Next tape exchanges are planned for today (#10 to #11), and 2/25 (#11 to #12).  A new table will be uploaded in order to modify the PROMISS video recording rate.  No on-orbit recordings are foreseen for 2/25 and 2/26 due to the EVA activities.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Renal Stone (RS):   Last in-flight data collection session will occur in April.  Thanks to the crew for their continued support with taking their daily pills at dinner time.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Pre-Increment requirements have been completed.  A second BBT (Beacon & Beacon Tester) session will be scheduled in US Lab.

Foot/Ground Reaction Forces during Space Flight (FOOT):   A second downlink of the Foot data has been requested.  Once the data is on the ground and the PI (principal investigator) has reviewed the file, the ground will get back to Foale with the results.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Looking forward to the next set of FDI Tissue Culture runs.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new. 

Earth Science Toward Exploration Research (ESTER):  Looking forward to the upcoming sessions.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   CGBA continues to run nominally at 20C.

Miscible Fluids in Microgravity (MFMG):    Nothing new.

Educational Payload Operations (EPO):  The education demonstrations continue to be excellent.  The video will be used in a variety of ways.  The ground is looking forward to upcoming activities.

Crew Earth Observations (CEO):   A sharp ISS/CEO image of the crescent Moon and the upper atmosphere of the Earth, shot back in December, will be published on Earth Observatory this weekend with a short article highlighting a NRL (Naval Research Laboratory) study reporting a decrease in the density of the thermosphere.  With more targets for the crew to see in the Northern Hemisphere, their volume of imagery has increased dramatically the past couple of weeks, and the ground has quite a backlog to review.  Researchers have cataloged several recent sessions over their target area of the Bahamas and found good imagery, but focus is still a bit soft with the long lens views.  The crew responded very well to the uploaded dynamic event targets of ice rafts on Lake Superior and Ice Floes on James Bay.  They did a near-perfect job of documenting a major point source for dust northeast of Lake Chad.  A quick look at their 2/16 images of dynamic dust and sandstorm events that they reported over northern Africa and SW Asia offers some incredible views that will be studied more carefully soon.

Today’s optional CEO targets, with targets in Asia excluded due to the current sleep cycle, were Paleo-river patterns, SW Libya (good pass over the northeast part of the site: suggested were views to nadir and S to capture most of this site in general views — for future homing in on the meandering ancient subglacial river courses), Dust event, Chad (Dynamic event.  Starting about two days ago, a new major dust event seems to be setting up in the Sahel.  Crew was to look right for thick dust in the major depression of central Chad.  The last Saharan dust event exported dust from Chad to South America), Dust event, Mali (Dynamic event.  Looking left for a major air mass boundary:  clearer [moister] air to the N and dustier [drier] air to the S [visibilities only one mile in Niamey, the capital city]), Lagos, Nigeria (nadir pass over this coastal city), Johannesburg, South Africa (nadir pass.  Two views, one slightly left of nadir and one slightly right, will capture this major city.  Johannesburg’s gold mines and industry used to attract workers from as far north as the Congo.  Now the economic pull attracts traders from Nigeria and North Africa), Midway Island, Pacific (400mm-lens.  Nadir pass), Kure Atoll, Pacific (400mm-lens.  Nadir pass.  This is the most northerly coral atoll in the world), and Tuamotu Archipelago (skies should have been clear over the southern half of the island chain.  Detailed images of the coral atolls were requested).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

ISS Orbit  (as of this morning, 11:15am EST [= epoch]):

  • Mean altitude — 366.5 km
  • Apogee — 370.9km
  • Perigee — 362.1 km
  • Period — 91.88 min.
  • Inclination (to Equator) —  51.627 deg
  • Eccentricity — 0.0006477
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 90 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 30003

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.